OpenStudy (anonymous):

The inverse of g(x) = -1/x is

OpenStudy (johnweldon1993):

Inverse means switch the 'x' and the 'y'...and then solve for 'y' again So here \[\large y = \frac{-1}{x}\] will become \[\large x = \frac{-1}{y}\] Now how would you solve for 'y' again? *hint...don't over think it :D*

OpenStudy (anonymous):

I don't know! Make y equal 1?

OpenStudy (johnweldon1993):

No...in order to solve for 'y' again...you just use algebra to rearrange this so y = something \[\large x = \frac{-1}{y}\] Multiply both sides by 'y' \[\large x\times y = \frac{-1}{\cancel{y}}\times \cancel{y}\] So we have \[\large xy = -1\] And now to solve for 'y' divide both sides by 'x' \[\large \frac{\cancel{x}y}{\cancel{x}} = \frac{-1}{x}\] Which leaves us with \[\large y = \frac{-1}{x}\] Showing us...that y = -1/x has no inverse!

OpenStudy (anonymous):

1/x -1/x x -x so the answer is -1/x